OpenStudy (anonymous):

Which is the exact value of the expression the square root of 32. − the square root of 50. + the square root of 128.? 2the square root of 7. 7the square root of 2. 22the square root of 2. 2the square root of 55.

OpenStudy (anonymous):

\[\sqrt{32}-\sqrt{50}+\sqrt{128}\]

OpenStudy (anonymous):

so u needto simplify

OpenStudy (anonymous):

\[\sqrt{32}=\sqrt{16}*\sqrt{2}=4\sqrt{2}\] \[-\sqrt{50}=-\sqrt{25}*\sqrt{2}=-5\sqrt{2}\] and \[\sqrt{128}=\sqrt{64}*\sqrt{2}=8\sqrt{2}\] so rewrite ur equation substituting the values and u get: \[4\sqrt{2}-5\sqrt{2}+8\sqrt{2}=(4-5+8)\sqrt{2}\]

OpenStudy (anonymous):

simplify whats in the parenthasis and u will get us answer